MODS Technical Observing, prep for Partner Science:

Daily, during science blocks:

  • Check MODS status: from any partner account run mods1 status. You should find all systems running and owned
by "mods" (If they're not owned by "mods", that is OK, but it is good practice to start these at the console/KVM in CRB from "mods1data", in which case they will be owned by "mods"). If it is the start of a partner block, the User Interface probably will not be running; the partner will start it from their account and the owner will be the partner username.
lbto@obs3:14 % mods1 status
MODS1 Status at 2012 Nov 02 17:28:26
Instrument Server:
  Service   State      Owner
  --------------------------------------
  mmcService   Running      mods
  agwServer   Running      mods
  redIMCS   Running      mods
  blueIMCS   Running      mods
  lbttcs   Stopped
  modsUI   Stopped
  --------------------------------------

Data Server:
  Service   State      Owner
  --------------------------------------
  isis      Running      mods
  Red Caliban   Running      mods
  Blue Caliban   Running      mods
  modsenv   Running      mods
  --------------------------------------
  • If the IIF (TCS) is stopped and restarted, lbttcs will not be running and need to be manually restarted with mods1 start lbttcs
  • Quit and restart the CCD control programs (the ICs): (Restarting the ICs is necessary to clear the IC FITS header cache. There is a memory leak of some kind which has the result that two DATE-OBS keyword/value pairs, one stale, can be written to the data headers. The latter, stale one, is read by the archiving software which copies the file to the incorrect UTdate directory of the Repository)
    • Open the MODS KVM in Computer Room B, use power button to turn it on if it is off
    • Scroll-Lock Scroll-Lock to get the main KVM page
    • Select M1.RC and click connect
    • Type quit (nothing happens for ~2 seconds, then you get a C:\MODS> prompt at upper left)
    • Type IC. Screen will clear, wait for the status table at lower left to be populated
    • Repeat for M1.BC

Starting the CCD control programs is discussed in more detail within the MODS1 Instrument Startup Procedure. See this page if more details are needed.

  • start the User Interface if it is not open already: mods1 start gui
  • wake up MODS: _execMODS
  • modsDisp if it is not running already
  • adjust Setup information: set name of Support and TelOps and Get Date
  • Confirm Calibration mode is set and next filenames are correct
  • execMODS ~MODSeng/Support/sieveSnap.pro
  • verify data taken, displayed, saved to archive
    • central pixel should be
  • put MODS into Observing Mode from GUI
  • ask OSA to do their test preset

Before each science block (afternoon and science prep night tasks)

  • Run Daily Checks, above.
  • I like to run, in addition to sieveSnap.pro, a script which takes dual blue/red images through the imaging mask (execMODS /home/LBTO/SciPrep/MODS/imSnap.pro)
  • Organize/Purge MODS data disk
    • must do this as user mods on mods1data, so ssh mods@mods1data
    • periodically make UTdate subdirectories under /lhome/data and move data into these:
      • cd /lhome/data
      • mkdir UTdate (e.g. mkdir 20110929)
      • mv mods1?.20110929.*.fits 20110929/
    • Total disk space on current system is about 100 Gb. A good nights data, with calibrations, will use 6-8 Gb. When the disk becomes full, you can delete older UTdate directories.
      • du -sh (lists how much space is used by each subdirectory, in a human-readable form.
      • To delete an entire directory rm -rf 20110616
  • Observe a spectrophotometric standard in all configurations
    • Calibration scripts in ~MODSeng/modsScripts/modsSpecPhot
    • If it is not photometric, try to remember to note this in the title (copy and edit the script) and ask the OSA to make sure to note it in their log.
    • One of the primary spectrophotometric standards would be best, since these have better coverage in the red.
    • Include table of name, RA, DEC of available standards here
    • This should take ~30 minutes - the stars have a range of brightnesses and, thus, exposure times (Note: we may make some versions to take only 1 image per configuration, so with nimgs 1).
  • Check that offsets are accurate (scripts in /home/LBTO/SciPrep/MODS/AlignCheck/)
    • 5-dice with detxy:
      • execMODS dice5_xy_02+31.img
      • execMODS dice5_relxy_02+31.img
    • 5-dice with radec:
      • execMODS dice5_relrd_02+31.img
      • execMODS dice5_rd_02+31.img
    • Blind offsets (script not prepared yet but dice5_relrd+02+31.img is a similar check).

Investigating known issues (tasks not for each science prep night but to investigate issues which have arisen)

  • Stray light (Scripts are in /home/lbto/SciPrep/MODS/StrayLight)
    • 0751 field showed very strong stray light pattern (300 counts over background) and is the best field for this investigation. It rises around 01 MST (08 UT). 0305 field also showed a strong pattern (100 counts) and is up earlier. It would be preferable to observe 0751 if possible.
    • Color dependence: SL0751dual_colors.img or SL0305dual_colors.img
    • Depedence of pattern on offset of bright star from field center: SL0751dual_cardoff.img or SL0305dual_cardoff.img
    • Other scripts to make large (30 arcmin) offsets or smaller (1 arcmin) offsets are SLxxxxdual_bigoff.img and SLxxxxdual_offsets.img, respectively.
    • Scripts to observe the fields in red-only or blue-only mode, using all red or all blue filters, are also in this directory.

Annual tests, restart after summer shutdown:

Preparatory to Day or Nighttime Tests (to be done by ISS with InstSp) :
  • Check that glycol is flowing and check CCD temperatures (just trying to start the minimum necessary to confirm that glycol is flowing and the CCDs are cold)
    • To check temperatures, first check status and run the isis and env services:
      • mods1status --- check mods status
      • mods1 start isis --- start message service
      • mods1 start env --- start environmental logging
        • the env logfile will show the supply/return glycol temperatures and pressures. Also the air temperatures in the instrument. But it will not report dewar
      • Confirm that glycol is flowing by looking at the temperatures in the logfile and asking the telescope manager on duty.
    • To check dewar pressure and temperature, run the command "modsTemps 1 blue" or "modsTemps 1 red" (NB: I think the "IC"s additionally need to be turned on).
    • Instrument startup (follow "MODS1 Instrument Startup Procedure" on the TWiki at http://wiki.lbto.org/twiki/bin/view/Instrumentation/MODSStartUp, starting from the "Instrument Control Software Startup").
    • The 4 MODS1 computers (mods1, mods1data, M1.BC and M1.RC) should already be powered up, hence the advice to start from "Instrument Control Software Startup". If these computers are not running, then start from the top of this Instrument Startup Procedure TWiki page).
  • After all MODS1 services are running, execute the script to do a Cold Start. "execMODS modsColdStart.pro". A cold start homes and resets all of the mechanisms.

Daytime:
  • run a sieveSnap, imSnap (imaging mask instead of sieveMask, in /home/LBTO/SciPrep/MODS/) and a set of biases. This is a basic health check.
  • remeasure IMCS ttf zeropoints. Procedure on Twiki under Instruments->MODS. Compare with values on mods computer.
  • run through same set of calibrations as before dismounting MODS in July.
    • With telescope at zenith, and dome dark:
      • sieveSnap.pro
      • grlamps.cal
      • grflats.cal
      • prlamps.cal
      • prflats.cal

Nighttime:

The first ones come from the MODS1 AGw Commissioning report - sections 5.1.2 and 5.1.3 of "On-Sky Calibration and Tests". As far as the remainder of section 5, I do not think these steps must be repeated since the MODS1 AGw is part of MODS1. But I think that obtaining at least a set of 5 5-dice dithers on a dense star field would be a good sanity check on the accuracy and repeatability of offsetting.
  • Check hotspots: guidecam_center and guidecam_hotspot (no change expected - not essential)
  • Re-measure rotator center. Procedure in AGw Commissioning Report, section 5.1.2. (should be done)
  • Re-measure rotator zero point (LEFTZEROPOINT). Procedure in AGw Commissioning Report, section 5.1.3. This involves, after collimation, sending a track preset, and then doing a "slitscan", i.e. making a series of Y-offsets (detxy 0 dy). Finally take a 120-sec image of the 0.6" slit. (should be done)
    • Scripts are in lbto@obs3:/home/lbto/SciPrep/MODS/SlitScan and are StoneL1_slit.img, StoneOgs1224_slit.img and StoneOgs1344_slit.img. Only one needs to be run. The Stone L is up at the start of the night (RA~19, DEC=0), and O (RA~21:45, DEC=0) rises soon after. StoneOgs1224 uses a bright guide star
but in the MODS field. StoneO1344 uses a guide star off the field. Interesting to compare collimation, but just do the StoneO1344 if there is a choice.
  • 5-dice (5x) on Stone field. Should take about 40 minutes. (should be done).
    • Scripts are in lbto@obs3:/home/lbto/SciPrep/MODS/AlignCheck/.
      • StoneO_1136_act0_5dice5.img: RADEC, absolute offsets, 5x dice 5, 10" offsets
      • StoneO_1136_act0_5dicexy5.img: DETXY, absolute offsets, 5x dice 5, 10" offsets
      • StoneO_1136_act0_5dicerelrd.img: RADEC, relative offsets, 1x dice 5, 10" separation
      • StoneO_1136_act0_5dicerelxy.img: DETXY, relative offsets, 1x dice 5, 10" separation
  • Spectrophotometric standard in all configurations (please label as non-phot if not photometric, and try to obtain a second set when photometric). Should take about 30 minutes on average, but depends on star chosen. (should be done)
    • Check positioning of standard star in LS60x5 slit after "modsAlign -r" acquisition (should be done)

Additional September 2014 restart tasks:

  • in afternoon, run imflats_edit.cal in ~lbto/SciPrep/MODS/modsCalibs_edit (or something like that)
  • MODS + LBC-R pseudo-monocular observation with MODS dithering and LBC-R guiding
    • Any conditions OK for functionality test
    • Scripts in ~lbto/SciPrep/MODS/PseudoMono
      • PDS456 good until 11pm
      • FF2131 for later (MODS and LBC-R I-BESSEL) need to observe Feige110 also in 0.6" slit..
  • MOS Alignment
    • Conditions: Seeing should be less than ~1.5-2".
    • Scripts in ~lbto/SciPrep/MODS/MOSAlign
      • acqMODS for one of the modsplnk fields - masks are loaded
      • run either of the img scripts to offset rot/x/y - choice will depend on time, StoneO is up until xx, 02 is up until xx.
  • Details of guide camera rotator center measurement:
    • Conditions: any
    • Ask the OSA to sent an active on-axis preset to a BSxxxx star and collimate.
    • Once collimated:
      • Stop Guiding
      • dIE 5 (offset telescope)
      • readGuideCam right -e 1000 (dummy exposure because there are usually problems on the first exposure after stopping guiding)
      • Move rotator to 0, then slew to 360
        • readGuideCam right -e 90000
      • can repeat 1-2 times and measure the center on the guide camera image.
  • If photometric:
    • Extinction curve: Observe BD28 from zenith to setting in dual grating mode.

Longer Term Improvements

  • Measure Mt.Graham mean extinction coefficients for optical spectroscopy. (Note, this was done on
    • Needs photometric weather!
    • Preference for primary standards because of sampling at extreme wavelengths
    • Need range of air masses (at least 1.25 - 2.20, our 25d limit is 2.367)
Topic revision: r26 - 11 Sep 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback